OpenStudy (anonymous):

((PLEASE HELP FAST!!!) Solve each one with a graphing calculator /which I have no idea how to use/ 6x-3 =52 10x+4 = 1000 7x = 124 2x-4 + 10 = 22 3(5)2x-3 = 6

OpenStudy (oaktree):

Oh my. You should really not be using a graphing calculator for these. I'm going to assume you have a TI-Nspire, which is really the only calculator which does these things. In that case, look up the user's manual - it should have instructions.

OpenStudy (anonymous):

A graphing calculator is the last thing I want to use, but unfortunately for me, my teacher requires me to "solve by graphing" :(

OpenStudy (oaktree):

Oh, god. You're kidding me, right? OK, what calculator do you have?

OpenStudy (oaktree):

What does it look like? Does it have just numbers, or does it also have separate keys for the letters?

OpenStudy (anonymous):

\[6x-3 =52\\6x=49\\x=\frac{49}{6}\]\

OpenStudy (anonymous):

Are you not allowed to solve it algebraically? Graphing is probably the most inaccurate way to solve unless it will give you a table of values.

OpenStudy (anonymous):

Nope, sadly... Only graphing allowed :(

OpenStudy (anonymous):

\[10x+4 = 1000\\10x=996\\x=\frac{996}{10}=99.6\]

OpenStudy (anonymous):

https://www.wolframalpha.com/input/?i=7x+%3D+124

OpenStudy (anonymous):

\[7x=124\] \[x=124\div 7=17.714\]

OpenStudy (anonymous):

That helps a lot; thanks :)

OpenStudy (anonymous):

what kind of a nimrod math teacher wants you to solve simple linear equations using a graphing calculator????

OpenStudy (anonymous):

Haha, this teacher always seems to make things harder than they need to be. She doesn't even let me use the FOIL method on binomials

OpenStudy (anonymous):

\[2x-4 + 10 = 22\\2x+6=22\\2x=28\\x=28\div2 = 14\]